There are 7 books numbered 1–7 on the summer reading list. Peter randomly
chooses 2 books. What is the probability that Peter chooses books numbered 1 and 2?

Answer by stanbon(75887)

There are 7 books numbered 1–7 on the summer reading list. Peter randomly
chooses 2 books. What is the probability that Peter chooses books numbered 1 and 2?
---------------------------
Books 1 and 2 form one group.
There are 7C2 = (7*6)/(1*2) = 21 groups of two
------------------------------------------------
P(choosing the 1,2 group) = 1/21
===================================
Cheers,
Stan H.